Please solve the following equation by factors (Please do not use the quadratic formula)

2x squared -7x=15

Answer by ewatrrr(24785)


Hi,
2x^2 - 7x = 15
2x^2 - 7x -15 = 0
factoring
(2x + ?)(x - ?) will need to be the format of the factoring
(2x +3)(x -5) = 0 works as 3x - 10x = -7x
(2x +3)=0
x = -3/2
(x -5)=0
x = 5